OpenStudy (anonymous):

Find an example of a field F and elements a and b from some extension field such that F(a,b) ≠ F(a), F(a,b) ≠ F(b), and [F(a,b):F] < [F(a):F][F(b):F

OpenStudy (kinggeorge):

So the first the I would try to do, is find a degree 4 extension over some field (generated by a single element). Can you give me some extension/field that satisfies this?

OpenStudy (kinggeorge):

Well, in any case, the idea is as follows. Find fields \(\mathbb{F}\subseteq\mathbb{F}(\alpha),\mathbb{F}(\beta)\subseteq\mathbb{F}(\alpha,\beta)\) such that \([\mathbb{F}(\alpha,\beta):\mathbb{F}]=8\), and \([\mathbb{F}(\alpha):\mathbb{F}]=[\mathbb{F}(\beta):\mathbb{F}]=4.\) So I guess it might be more beneficial to think of a degree 8 extension first.

OpenStudy (kinggeorge):

I might suggest considering the extension\[\mathbb{Q}(\sqrt[4]2,i)/\mathbb{Q}.\]You will of course need to write it slightly differently, since \(\mathbb{Q}(i)\) has degree 2 over \(\mathbb{Q}\) and not 4.

OpenStudy (anonymous):

I'm sorry I stepped away from my computer! I don't think I'm following your choice for an extension field

OpenStudy (kinggeorge):

Did you understand why we might want a degree 8 extension?

OpenStudy (anonymous):

because it will be a higher degree than the field F?

OpenStudy (anonymous):

oh wait 8<4x4=16?

OpenStudy (kinggeorge):

Right. And I'm pretty sure that 8 is the smallest degree where we could have such a construction. Now we just have to be careful about what degree 8 extension we choose. Having the benefit of experience, the extension I gave above is an extension that will work - if you're careful about how you choose your generators.

OpenStudy (anonymous):

so a and b would be the generators? (I have no experience like you haha, so this is completely new to me)

OpenStudy (kinggeorge):

\(a,b\) would be generators. But the way I've written it, the extension doesn't quite meet our extensions. Can you think of another pair \(a,b\) such that\[\mathbb{Q}(a,b)=\mathbb{Q}(\sqrt[4]2,i)?\]

OpenStudy (anonymous):

well you said above that only Q(i) didn't work....so could you just keep a and make b=i^2? Would that give you the right degree?

OpenStudy (kinggeorge):

That would be a bad idea since \(i^2=-1\in\mathbb{Q}\). What if you tried a combination of \(\sqrt[4]2\) and \(i\)?

OpenStudy (anonymous):

\[\mathbb{Q} \left( a,b \right)= \mathbb{Q} \left( \sqrt[4]{2},i \right)\]

OpenStudy (anonymous):

so i times 4th root of 2?

OpenStudy (kinggeorge):

Right. If we consider \(\mathbb{Q}(\sqrt[4]2,i\sqrt[4]2)\) instead, we can see that \(\mathbb{Q}(\sqrt[4]2,i\sqrt[4]2)=\mathbb{Q}(\sqrt[4]2,i)\). But also, \[[\mathbb{Q}(\sqrt[4]2):\mathbb{Q}]=[\mathbb{Q}(i\sqrt[4]2):\mathbb{Q}]=4.\]Do you see this is?

OpenStudy (kinggeorge):

*Do you see why this is?

OpenStudy (anonymous):

I see the first part yes, but how does Q(i)= Q(i(4th root of 2))?

OpenStudy (kinggeorge):

That's not true. In fact, we don't want that to be true.

OpenStudy (anonymous):

then you have lost me

OpenStudy (kinggeorge):

There are a bunch of field extensions between \(\mathbb{Q}(\sqrt[4]2,i\sqrt[4]2)\) and \(\mathbb{Q}\) (exactly 8 if you're curious). One of those intermediate extensions is \(\mathbb{Q}(\sqrt[4]2)\). Another is \(\mathbb{Q}(i)\) and a third one is \(\mathbb{Q}(i\sqrt[4]2)\). All of these three extensions are not equal to one another. However, we do have the property that\[\mathbb{Q}(\sqrt[4]2,i)=\mathbb{Q}(\sqrt[4]2,i\sqrt[4]2).\]

OpenStudy (anonymous):

so my generators are a=4th root of 2 and b=i times 4th root of two? and that should give 8<16=4*4?

OpenStudy (kinggeorge):

Moreover, with regards to your problem, \[[\mathbb{Q}(\sqrt[4]2,):\mathbb{Q}]=[\mathbb{Q}(i\sqrt[4]2):\mathbb{Q}]=4\neq2=[\mathbb{Q}(i):\mathbb{Q}].\]So we see that\[[\mathbb{Q}(\sqrt[4]2,i\sqrt[4]2):\mathbb{Q}]=8\]while\[[\mathbb{Q}(\sqrt[4]2):\mathbb{Q}][\mathbb{Q}(i\sqrt[4]2):\mathbb{Q}]=4^2=16\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (kinggeorge):

You're welcome.